OpenStudy (anonymous):

can someone help?use factoring to find the solutions of 2x^2=-7x+30 answer choices.... a. -6 and 5 b. -5 and 6 c. -6 and 5/2 d. -2/5 and 6

OpenStudy (anonymous):

2x^2=-7x+30 2x^2+7x-30=0 (2x-5)(x+6)=0 2x-5=0 or x+6=0 x=5/2 or -6